In 2hours 40minutes a car travels 160 kilometers. at what average speed is the car traveling
Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

to find km/minute, divide the kms by the minutes.
-----

Answer by ikleyn(52817)

.

HINT.   2 hours and 40 minutes = 60*2 + 40 minutes = 120 + 40 minutes = 160 minutes.